OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!! A family refers to a group of two or more people related by birth, marriage, or adoption who reside together. In 2000 in country A the average family net worth was 360,000, and there were about 9.7x10^7 families. Calculate the total family net worth in country A in 2000.

OpenStudy (anonymous):

Convert 360,000 into scientific notation as 3.6x10^5 The total net worth = 3.6x10^5 x 9.7x10^7 = 34.92x10^12 = 3.492x10^13